An oxygen-16 ion with a mass of 2.66 × 10−26 kg travels at 7.33 × 106 m/s perpendicular to a 7.23 – T magnetic field, which makes it move in a circular arc with a 0.652-m radius. What positive charge (in x10-19) is on the ion? 


answer to 2 decimal places


1
Expert's answer
2021-05-26T12:55:33-0400

"F=ma=qvB \space and \space a =\frac{v^2}{r}"

"qvB=\frac{mv^2}{r}"

"q=\frac{mv}{rB}=\frac{(2.66\times 10^{-26}\times 7.33\times 10^6)}{(0.652\times 7.23)}=4.14\times10^{-20}C"